kierrakeithhill:

When Bella pushes a box with a mass of 5.25 kilograms with a force of 15.75 newtons, it accelerates at a rate of 2.5 meters/second2. What is the force due to friction?

foxey3:

According to the Newton's second law, the net force acting on the body is equal to the product of mass and the acceleration in the body. Let f be the friction force, F be the applied force. F - f = m a 15.25 - f = 5.25 x 2.5 f = 15.25 - 13.125 f = 2.125 N Thus, the friction force is 2.125 N.

kierrakeithhill:

thankyou but the closest option to yours is 2.60newtons